OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 6 x – y = 3

OpenStudy (anonymous):

x=2 y=-1

OpenStudy (nurali):

4x + 2y = 6 x – y = 3-----(1) multiply '1' by 2 and then add. 4x + 2y = 6 2(x – y = 3) 4x + 2y = 6 2x – 2y = 6 ------------------ 6x = 12 x = 2 put 'x' in equation '1' 2-y=3 y=-1

OpenStudy (anonymous):

use matrices?

OpenStudy (anonymous):